OpenStudy (anonymous):

will fan and medal!!! okay I am working on a conic sections problem... "what are the foci of the elipse? Graph the ellipse. I have graphed the ellipse but I am having trouble solving for the foci.

OpenStudy (anonymous):

Here is the equation \[\frac{ x^2 }{ 49 }+ \frac{ y^2 }{ 64 }=1\]